A man can row his boat 10 kph in still water. In the same length of time that it took him to row 20 km downstream, he was able to row only half the distance upstream. What was the rate of the current?
Answer by richwmiller(17219)

r*t=d
(10-c)*t=10
10t-c*t=10
(10+c)*t=20
10+c*t=20
10t-ct=10
add
20t=30
t=30/20=1.5
t=1.5 hr
(10+c)*t=20
10t+c*t=20
10*1.5+c*1.5=20
15+c*1.5=20
c*1.5=5
c=5/1.5
c=10/3
c=3 1/3 kph current
check
10t+c*t=20
10*1.5+10/3*3/2=20
15+5=20
10t-ct=10
10(3/2)-10/3*3/2=10
15-5=10
ok